# PianoFlow: Music-Aware Streaming Piano Motion Generation with Bimanual Coordination

### 14 Apr 2026 v1

@@ 2.1. Audio-Driven Motion Generation
2.2. Piano Hand Motion Generation
@@ 3.1. Music-Aware Wrist Trajectory Generation
Given an input audio sequence, this stage predicts bimanual wrist trajectories $\varphi=[\varphi^{L},\varphi^{R}]\in\mathbb{R}^{N\times 6}$, where $N$ denotes the sequence length. To explicitly ground these trajectories in musical semantics, a cross-modal distillation scheme transfers structured harmonic priors from a multimodal (MIDI and audio) teacher to an audio-only student. Considering the distinct behaviors of each hand, we model their trajectories using independent networks with identical architectures but unshared parameters. This decoupled design prevents premature coupling, allowing each branch to capture hand-specific kinematic priors.
@@
To capture the hierarchical nature of musical intervals within each frame $n$, we model the local harmonic structure along the pitch dimension using a cascaded convolutional design: an octave-aware kernel (size $13$) followed by an interval-aware kernel (size $5$). To retain the spatial keyboard topology, we explicitly inject an absolute pitch positional encoding $\mathbf{P}_{\text{pitch}}\in\mathbb{R}^{88\times C}$. This yields a pitch-wise structural representation:
@@ 3.2. Bimanual Coordinated Gesture Generation
3.3. Streaming Inference via Autoregressive Flow Continuation
@@
where $\mathbf{m}$ decays from $1$ to 0 across the overlap region. This mechanism enforces boundary consistency while smoothing the transition into free generation. Executing this update at every solver iteration ensures that each new chunk integrates seamlessly with the prior gesture trajectory, maintaining long-term temporal coherence without inducing additional latency.
@@ 4.1. Experiments Setting
We evaluate the synthesized motions across three dimensions, including distributional fidelity, kinematic accuracy, and computational efficiency. This evaluation incorporates five base metrics (Gan et al., 2024) along with two extended metrics. (1) Fréchet Inception Distance (FID) quantifies the global distributional discrepancy between synthesized motions and ground-truth data within a pre-trained latent space. (2) Fréchet Gesture Distance (FGD) assesses hand-specific pose distributions by calculating the Fréchet distance within the MANO (Romero et al., 2017) parameter or joint coordinate space. (3) Wasserstein Gesture Distance (WGD) evaluates distributional alignment via optimal transport, calculating the Earth Mover’s Distance (EMD) following PCA and Gaussian Mixture Model (GMM) fitting. (4) Position Distance (PD) quantifies wrist spatial accuracy via the mean Euclidean distance between predicted and ground-truth wrist coordinates across the sequence. (5) Smoothness evaluates physical plausibility by measuring acceleration discrepancies relative to the ground truth to ensure temporal stability. (6) Real-Time Factor (RTF) measures computational efficiency, where an $\text{RTF}<1.0$ indicates real-time generation capability. (7) Frequency Domain Error (FDE) evaluates preservation of high-frequency kinematic details and penalizes over-smoothing artifacts from interpolation-based fusion. We compute amplitude spectra $S(f)$ and $\hat{S}(f)$ for the sequences via Fast Fourier Transform. Defining $\mathcal{H}$ as high-frequency bins exceeding $5$ Hz, FDE is the Mean Absolute Error (MAE) between spectral components to ensure precise motion alignment:

### 15 Apr 2026 v2

@@ 2.1. Audio-Driven Motion Generation
$2.$ Piano Hand Motion Generation
@@ 3.1. Music-Aware Wrist Trajectory Generation
Given an input audio sequence, this stage predicts bimanual wrist trajectories $\varphi=[\varphi^{L},\varphi^{R}]\in\mathbb{R}^{N\times 6}$, where NN denotes the sequence length. To explicitly ground these trajectories in musical semantics, a cross-modal distillation scheme transfers structured harmonic priors from a multimodal (MIDI and audio) teacher to an audio-only student. Considering the distinct behaviors of each hand, we model their trajectories using independent networks with identical architectures but unshared parameters. This decoupled design prevents premature coupling, allowing each branch to capture hand-specific kinematic priors.
@@
To capture the hierarchical nature of musical intervals within each frame $n$, we model the local harmonic structure along the pitch dimension using a cascaded convolutional design: an octave-aware kernel (size $13$) followed by an interval-aware kernel (size 55). To retain the spatial keyboard topology, we explicitly inject an absolute pitch positional encoding $\mathbf{P}_{\text{pitch}}\in\mathbb{R}^{88\times C}$. This yields a pitch-wise structural representation:
@@ 3.2. Bimanual Coordinated Gesture Generation
$3.$ Streaming Inference via Autoregressive Flow Continuation
@@
where $\mathbf{m}$ decays from 11 to 0 across the overlap region. This mechanism enforces boundary consistency while smoothing the transition into free generation. Executing this update at every solver iteration ensures that each new chunk integrates seamlessly with the prior gesture trajectory, maintaining long-term temporal coherence without inducing additional latency.
@@ 4.1. Experiments Setting
We evaluate the synthesized motions across three dimensions, including distributional fidelity, kinematic accuracy, and computational efficiency. This evaluation incorporates five base metrics (Gan et al., 2024) along with two extended metrics. (1) Fréchet Inception Distance (FID) quantifies the global distributional discrepancy between synthesized motions and ground-truth data within a pre-trained latent space. (2) Fréchet Gesture Distance (FGD) assesses hand-specific pose distributions by calculating the Fréchet distance within the MANO (Romero et al., 2017) parameter or joint coordinate space. (3) Wasserstein Gesture Distance (WGD) evaluates distributional alignment via optimal transport, calculating the Earth Mover’s Distance (EMD) following PCA and Gaussian Mixture Model (GMM) fitting. (4) Position Distance (PD) quantifies wrist spatial accuracy via the mean Euclidean distance between predicted and ground-truth wrist coordinates across the sequence. (5) Smoothness evaluates physical plausibility by measuring acceleration discrepancies relative to the ground truth to ensure temporal stability. (6) Real-Time Factor (RTF) measures computational efficiency, where an $\text{RTF}<1.0$ indicates real-time generation capability. (7) Frequency Domain Error (FDE) evaluates preservation of high-frequency kinematic details and penalizes over-smoothing artifacts from interpolation-based fusion. We compute amplitude spectra $S(f)$ and $\hat{S}(f)$ for the sequences via Fast Fourier Transform. Defining $\mathcal{H}$ as high-frequency bins exceeding 55 Hz, FDE is the Mean Absolute Error (MAE) between spectral components to ensure precise motion alignment: